Circular No. 8645

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005lt AND 2005lu Further to IAUC 8641, H. Khandrika, M. Baek, and W. Li report the

LOSS discoveries of two apparent supernovae on unfiltered KAIT images.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005lt Dec. 10.55 11^{h}42^{m}26.^{s}21 +20^{o}07′05′′.0 15.7 24′′.5 E, 5′′.4 S
2005lu Dec. 11.27 2 36 03.71 −17 15 50.0 17.1 1′′.8 E, 2′′.2 S

Additional magnitudes for SN 2005lt in MCG +03-30-51: May 29.20 UT, [19.0; Dec. 12.54, 15.7 (in poor conditions, but the new object was obviously detected). The offsets above for SN 2005lu are with respect to the nucleus of MCG –03-07-40, which is the western component of the interacting galaxy pair ESO 545-G38 (the eastern component is MCG –03-07-39). Additional magnitudes for 2005lu: Nov. 6.28 UT, [18.5; Dec. 14.13, 16.8.

SUPERNOVA 2005lv IN UGC 2964 T. Puckett and R. Gagliano report the discovery of an apparent su-

pernova (mag 18.7) on an unfiltered CCD frame taken with the Puckett Observatory 0.60-m automated supernova patrol telescope (cf. IAUC 8635) on Dec. 6.19 UT (and confirmed at mag 18.6 by Puckett on frames taken on Dec. 11.18 with a 0.60-m reflector). The new object is located at α = 4^{h}08^{m}56^{s}.47, δ = +27^{o}11′50′′.8 (equinox 2000.0), which is 13′′.5 west and 2′′.4 south of the center of UGC 2964. SN 2005lv was not present on images taken by Puckett on Nov. 23 (limiting mag ∼ 20.0).

V838 MONOCEROTIS D. K. Lynch, R. J. Rudy, S. Mazuk, and C. C. Venturini, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; R. B. Perry, Langley Research Center, NASA; and B. Walp, Lick Observatory, report 0.47- to 2.5-μm spectroscopy of V838 Mon (cf. IAUC 7785, 7812, 8221) using the Lick 3.0-m telescope (+ VNIRIS) on Nov. 13.42 UT. The object continues to develop spectrally and to fade. Since 2004 there has been a dramatic weakening, and in some cases disappearance, of the TiO and VO absorption bands, probably due to thinning of the line-of-sight density as the shell expands. The AlO bands also weakened, though only slightly. There were also a large number of molecular absorption bands in the visible. Approximate magnitudes for V838 Mon: V = 15.7, J = 7.8, H = 6.6, K = 5.8.

2005 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green